# Pixel column issue in the ATLAS Inner Tracker modules

### v2

@@ 3.2 Inconsistencies and analog dependencies
Very few CC cases can be narrowed down to bad pixel regions (4$\times$1 pixels) as shown in figure 5(a) or bad pixel cores (8$\times$8 pixels) as shown in figure 5(b). Once these smaller areas are disabled, the chip can function normally without disabling the CCs. This would drastically reduce the number of bad pixels overall on a chip compared to disabling entire core columns.
However, CC issue often shows inconsistent symptoms even within the same chip. Another example of the CC issue is shown in figure 5(b), the software diagnostics tool cannot find any bad core columns. The analog scan sometimes runs successfully but then bad pixel cores can show up in the occupancy maps.
@@
Figure 5: (a) A bad 4$\times$1 pixel region. (b) A bad 8$\times$8 pixel core. (c) Occupancy map of a module at 1650 e injection charge. (d) Occupancy map of the same module at 2600 e injection charge.

### v3

@@ 3.2 Inconsistencies and analog dependencies
Very few CC cases can be potentially narrowed down to bad pixel regions (4$\times$1 pixels) as shown in figure 5(a) or bad pixel cores (8$\times$8 pixels) as shown in figure 5(b). Once these smaller areas are disabled, the chip can function normally without disabling the CCs. This would drastically reduce the number of bad pixels overall on a chip compared to disabling entire core columns.
CC issue often shows inconsistent symptoms even within the same chip. One example is figure 5(b), where the bad pixel core was hard to find. While the software diagnostics tool cannot find any bad CCs in this chip, the analog scan often fails, indicating CC issues. Only when sometimes an analog scan runs successfully, this bad pixel core can be seen in the occupancy map.
@@
Figure 5: (a) Map from a pixel test highlighting a bad 4$\times$1 pixel region. (b) Occupancy map from an analog scan with a bad 8$\times$8 pixel core with mostly 0 hits. (c) Occupancy map of a module at 1650 e injection charge. (d) Occupancy map of the same module at 2600 e injection charge.